2. You are trying to decide how to most efficiently use your oven. You do not want the oven running at a high temperature when you are not baking, but you also do not want to waste a lot of time waiting for the oven to reach the desired baking temperature.
The instruction manual on the industrial oven suggests the oven temperature will increase by 45 degrees Fahrenheit per minute. When the oven is initially turned on, the temperature is 70 degrees Fahrenheit. What will the temperature of the oven be after 7 minutes? Write an expression and explain how you arrived at your answer.

Answer by Theo(13342)

the expression is:

45*7 + 70

oven starts at 70 degrees.

each minute it is turned on, it gets 45 degrees hotter.

45 * 7 = 315

315 + 70 = 385

general formula would be:

y = 45*x + 70

y is the temperature it reaches
x is the number of minutes it is on.

the domain of this function would be from 0 to the the maximum number of minutes that the temperature will increase.

the range of this function would be the maximum temperature that the oven will reach.

assuming the maximum temperature is 520 degrees, then:

to find the maximum number of minutes, simply replace y with 520 to get:

520 = 45*x + 70

subtract 70 from both sides of this equation to get:

450 = 45*x

divide both sides of this equation by 45 to get:

x = 450 / 45 = 10

the domain of this function is 0 <= x <= 10

the range of this function is the value of y when x = 0 and the value of y when x = 10

when x = 0, y = 70

when x = 10, y = 45*10 + 70 = 450 + 70 = 520.

this assumes that the oven will always be at 70 degrees as soon as it is turned on.